OpenStudy (anonymous):

what is the gcf for 24 and 30

OpenStudy (amistre64):

split them up and what do you get? another day older and deeper in debt

jimthompson5910 (jim_thompson5910):

Factor each into prime factors 24 = 2*2*2*3 30 = 2*3*5 Now highlight all the primes that both have in common.

jimthompson5910 (jim_thompson5910):

and take those primes and multiply them together to get the GCF

OpenStudy (anonymous):

24=2*2*2*3 30=2*3*5 Go from here.

OpenStudy (amistre64):

24 = 1,2,3,4, 6,8, 12, 24 30 = 1,2,3, 5,6, 10, 15, 30 ^ biggest factor in common is?

OpenStudy (anonymous):

factors of 24- 1,2,3,4,6,8,12,24 30- 1,2,3,5,6,10,15,30 look at what they have in common - 1,2,3,6, take the largest one, hence GREATEST COMMON factor 6 it's 6

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

You rock so much